# RUMOR: Reinforcement learning for Understanding a Model of the Real World for Navigation in Dynamic Environments

### v1

@@ 1 Introduction
In Section 2, we analyze the background and relate our work with the state of the art and in Section 3 we state the preliminaries needed for understanding the method. Section 4 and Section 5 present the approach taken. We provide comparisons of RUMOR with other method of the state-of-the-art and experiments with a real-ground robot in Section 6. Finally, the conclusions of the work are stated in Section 7.
@@
The code will be made open source upon acceptance.
2 Background
@@ 2.1 Motion planning in dynamic environments
Motion planners for static and continuous environments do not properly deal with dynamic obstacles, as they lead to collisions and suboptimal trajectories, as seen in some works as the benchmark proposed in [2]. Some traditional approaches for dynamic environments include forces to model the obstacle avoidance, such as artificial potential fields [3] or pedestrian forces to emulate the people motion [4]. They were later improved with Time Elastic Bands (TEB) [5] or multiple extensions to Social Force [6] [7].
A big group of works are velocity-space based. The Velocity Obstacle (VO), introduced in [8], refers to the set of velocities of the robot that could lead to a collision with an obstacle that has a certain velocity in the near-future, which should not be chosen. Based on the VO concept, the Dynamic Object Velocity Space (DOVS) is defined in [1] as the velocity-time space for differential-drive robots, which includes the unsafe robot velocities for all the obstacles and the time to collision for computing safe robot velocities in a time horizon. In that work, a planner based on strategies is also defined, the S-DOVS. Another work [9] uses the DOVS to discard velocities that lead to collisions from the action space of a tables-based reinforcement learning algorithm. However, unlike our work, it does not use the information of the DOVS to interpret the environment and just filters the unsafe actions available. In addition, the results presented in their work are very similar to S-DOVS, which does not require training time. Other extensions of VO use reciprocal collision assumptions, such as Reciprocal Velocity Obstacles (RVO) [10] or Optimal Reciprocal Collision Avoidance (ORCA) [11], which leads to failures in scenarios where there are non-collaborative obstacles.
@@ 2.2 Deep reinforcement learning planners
Deep Reinforcement learning [16] is a method used to learn to estimate the optimal policy that optimizes the cumulative reward obtained in an episode with a deep neural network. Many extensions have been proposed to this original algorithm. Some works have proven the best performances of the state-of-the-art in various problems, including Rainbow [17] for discrete action spaces, distributed reinforcement learning [18] or actor-critic methods [19].
Some works offer analysis of the importance of Deep Reinforcement Learning (DRL) in robot motion planning and the limitations of traditional planners in dynamic environments. Defining strategies for every situation that may be found in the real world or optimizing trajectories in very dynamic environments is intractable, so DRL may be used to efficiently solve the decision-making problem, which is complex and has many degrees of freedom. There are works [20] [2] [21] that implement and compare DRL algorithms with model-based ones, proving their performance.
The first agent-based DRL approach proposed for motion planning [22] uses a fully connected network to select velocities regarding the position and velocity of the closest surrounding obstacles. LSTM [23] layers were included in [24] [25] to account for multiple obstacles. SARL [26] models with attention the interactions among obstacles, achieving state-of-the-art performance. Recent approaches compare their results to it, and they usually focus on improving the structure of the network. For example, [27] uses a graph neural network or [28] an attention-based interaction graph. A recent approach [29] improves the performance of others with intrinsic rewards. The use of DRL algorithms in robot navigation has many important challenges, including the need of learning from a limited subset of possible scenarios, environmental and robot constraints or partial observability [30] [31]. The previous approaches do not face these problems. In our approach, instead of directly process the observations, we use the DOVS model, that represents the dynamism of the environment in the velocity space, as an input to the DRL algorithm, doing the decision-making process agnostic of the specific scenario.
@@
3 Preliminaries
3.1 Problem formulation
We consider the problem of robotic navigation in dynamic environments. A robot, whose position at time $t$ is $\mathbf{p_{t}}=(x_{t},y_{t},\theta_{t})\in\mathbb{R}^{3}$ and its velocity command $\mathbf{u}_{t}=(\omega_{t},v_{t})\in\mathbb{R}^{2}$ (angular and linear velocity), has to reach a goal $\mathbf{g}=(x_{g},y_{g})\in\mathbb{R}^{2}$ in the minimum possible time, while avoiding static and dynamic obstacles in the environment. The robot and the obstacles share the workspace $\mathcal{W}=\mathbb{R}^{2}$. The robot is represented with a disk with radius $r_{robot}$, and the dynamic obstacles are assumed to be circular disks with radius $\mathbf{r}_{obs}\in\mathbb{R}^{M}$, where $M$ is the number of dynamic obstacles. Let $\mathcal{R}_{t}(\mathbf{u})$ be the robot state at time $t$ after applying the command velocity $\mathbf{u}$; and $\mathscr{O}_{\forall,t}=\bigcup_{i=1}^{M}\mathscr{O}_{i,t}$ be the set of points occupied by the whole set of static and dynamic obstacles in the environment at time $t$, and $\mathscr{O}_{\forall}$ at any time. The robot maximum linear and angular velocities are $v_{max}$ and $\omega_{max}$, and its minimum linear and angular velocities are 0 m/s and $-\omega_{max}$, respectively. Equation 1 represents the unicycle model,
@@ 3.1 Problem formulation
and the the low-level forward model that relates the wheel velocities to the actions $\mathbf{u}_{t}$ for a differential-drive robot, is the following:
@@
where $r_{w}$ is the wheel radius, $\omega_{r,t}$ and $\omega_{l,t}$ the angular velocity of the right and left wheels, respectively, and $L_{w}$ the distance between wheels. This poses a linear relationship between the maximum $v_{t}$ a robot may take in a particular moment regarding its $\omega_{t}$:
@@
being $v_{max}$ and $\omega_{max}$ the maximum linear and angular velocities, respectively, and $\frac{v_{max}}{\omega_{max}}=\frac{L_{w}}{2}$. The robot has a maximum linear acceleration of $a_{max}$. Thus, the change in the linear velocity of the robot is also limited by its previous linear velocity and, due to Equation 2, its velocity limits:
@@
The restrictions of Equation 3 and Equation 4 are plotted in a graphical representation of the robot velocity space as the two black lines and the green rhombus, respectively, in Figure 2.
Figure 2: Graphical representation of the DOVS model for a differential-drive robot velocity space, with the kinodynamic constraints of Equation 3 and Equation 4 represented with two black lines (velocities greater than them are not reachable) and a green rhombus (velocities outside the rhombus are not reachable in one control period), respectively. The current robot velocity is the center of the rhombus. In this example, the robot velocity limits are $v_{max}=0.7$ m/s, $\omega_{max}=\pi$ rad/s and $a_{max}=0.3$ m/s². The dark area (DOV) includes unsafe velocities and white area (FV) corresponds to safe velocities.
@@ 3.2 Dynamic Object Velocity Space (DOVS)
The DOVS model, presented in [1], is used to extract the environment information in this work. It is a model that represents safe and unsafe velocity commands for a differential drive robot within a temporal horizon in a dynamic scenario. It extracts the information of the workspace directly in the robot velocity space $\mathcal{V}$ in every control period, using the time to collision between the robot and obstacles trajectories.
To build the model regarding the dynamic obstacle $i$, the robot size is reduced to a point and the obstacle is enlarged with the robot radius (the final collision areas are the same). The collision band $\mathcal{B}_{i}$ is computed as the area swept by the enlarged moving obstacle using its trajectory, which is assumed to be known or estimated from the sensor information. We consider circular trajectories as constant linear and angular velocities leads to them. We denote $\tau_{j}\equiv\tau_{j}(\omega_{j},v_{j})$ the circular trajectory $\tau_{j}$ produced by $\omega_{j}$ and $v_{j}$. Let $\mathcal{T}$ be the collection of different circular trajectories that collectively cover $\mathcal{W}$. For each circular trajectory $\tau_{j}\in\mathcal{T}$, the intersection points between the two lines limiting $\mathcal{B}_{i}$ and $\tau_{j}$ are $P_{1,i,j}(x_{1,i,j},y_{1,i,j})$ and $P_{2,i,j}(x_{2,i,j},y_{2,i,j})$. They are graphically represented in Figure 3(a) in a robocentric view for a few example trajectories.
@@
Figure 3: A robocentric view of the workspace (a) and a graphical representation of the DOVS (b) of a scenario with a robot navigating and facing the obstacle $i$ that follows a linear trajectory. In the robocentric view, the robot center is represented in red, trajectories $\tau_{j}$ sampled in blue circular lines, the obstacle augmented radius with a blue circle and the collision band $\mathcal{B}_{i}$ with green lines. The intersection points between $\tau_{j}$ and $\mathcal{B}_{i}$ are $P_{1,i,j}$ (right) and $P_{2,i,j}$ (left). In the DOVS, the maximum velocities to pass after the obstacle are represented with black dots ($\omega_{2,i,j},v_{2,i,j}$) and the minimum velocities to pass before it with purple dots ($\omega_{1,i,j},v_{1,i,j}$). The resulting unsafe velocities are represented in red.
The time $t_{2,i,j}$ that obstacle $i$ takes to reach $P_{2,i,j}$ and the time $t_{1,i,j}$ that it takes to stop being in contact to $P_{1,i,j}$ measure the times of collision between the robot and the obstacle. This means that $\tau_{j}$ will intersect $\mathscr{O}_{i,t}$ for $t\in[t_{1,i,j},t_{2,i,j}]$ (see [1] for more details about this computations). Therefore, the velocities that let the robot avoid collisions with obstacle $i$ within trajectory $j$ are those which are limited by the minimum and maximum velocities that allow the robot to pass before or after the obstacle. They are computed as:
@@
where $r_{j}$ is the radius of $\tau_{j}$ and $\theta_{k,j}$ the angular displacement on $\tau_{j}$ to reach $P_{k,i,j}$. Therefore, the set of velocities $\left\{(w,v)\in\mathcal{V}\mid v_{1,i,j}\leq v\leq v_{2,i,j}\;\land\;\omega=\frac{v}{r_{j}}\right\}$ lead to collision at a given time within the interval $t\in[t_{2,i,j},t_{1,i,j}]$. The Dynamic Object Velocity (DOV) for obstacle $i$ (red area in Figure 3(b)) and the set of trajectories $\mathcal{T}$ is defined as:
$\begin{split}DOV(\mathscr{O}_{i,t},\mathcal{T})=&\left\{\mathbf{u}_{j}=(\omega_{j},v_{j})\in\mathcal{V}\mid\exists\tau_{j}\in\mathcal{T},\;\right.\\ &\left.\exists t\in[0,T^{h}],\;\mathcal{R}_{t}(\mathbf{u}_{j})\cap\mathscr{O}_{i,t}\neq\emptyset\right\},\\ \end{split}$ (6)
where $r_{j}$ is the radius of the circular trajectory $\tau_{j}$ and $T^{h}$ a time horizon. The DOV may be combined for all obstacles in the environment:
$DOV(\mathscr{O}_{\forall,t},\mathcal{T})=\bigcup DOV(\mathscr{O}_{i,t},\mathcal{T})$ (7)
$DOV(\mathscr{O}_{\forall,t},\mathcal{T}$) intuitively represents the unsafe velocities that sometime lead to collision with the object. The Free Velocity (FV) is the complementary set of the DOV within $\mathcal{V}$, and represent the set of safe velocities that do not lead to collision with any obstacle.
$\begin{split}FV(\mathscr{O}_{\forall,t},\mathcal{T})=&\left\{\mathbf{u}_{j}=(\omega_{j},v_{j})\in\mathcal{V}\mid\mathbf{u}_{j}\notin DOV(\mathscr{O}_{\forall,t},\mathcal{T})\right\}\\ \end{split}$ (8)
Finally, the DOVS is defined as the whole set of velocities that comprise the DOV and the FV:
$DOVS(\mathscr{O}_{\forall,t},\mathcal{T})=\left\{DOV(\mathscr{O}_{\forall,t},\mathcal{T})\cup FV(\mathscr{O}_{\forall,t},\mathcal{T})\right\}$ (9)
The DOVS contains the DOV representing unsafe velocities leading to collision in a time horizon if they are kept and FV contains safe velocities. The model may be reactively computed in every control period, giving the robot updated information to navigate adapted to the precision of the information sensed. Velocities inside the DOV do not directly lead to a collision if the following commands lead to a free velocity. This is represented in the evolution of the DOVS in time.
In DOVS, the velocities reachable for a differential-drive robot are constrained by the Equation 3 and 4, as shown in Figure 2. Therefore, only the velocities accomplish the constraints of the previous equations, are reachable in the next control period from the current velocity. The whole model is used to learn and apply the robot velocities for a safe navigation, as explained in the following sections.
4 Reinforcement learning setup

### v2

@@ Abstract
©This paper has been accepted for publication at Robotics and Autonomous Systems. Please, when citing the paper, refer to the official manuscript with the following DOI: 10.1016/j.robot.2025.105020.
@@ 1 Introduction
In Section 2, we analyze the background and relate our work with the state of the art and in Section 3 we state the preliminaries needed for understanding the method. Section 4.2 and Section 4.3 present the approach taken. We provide comparisons of RUMOR with other method of the state-of-the-art and experiments with a real-ground robot in Section 5. Finally, the conclusions of the work are stated in Section 6.
@@
The code and videos are available in https://github.com/dmartinezbaselga/RUMOR.
2 Related work
In this section, we give a brief overview of the existing solutions for autonomous navigation in dynamic environments, with special emphasis in deep reinforcement learning planners. We refer the readers to recent surveys [2, 3, 4] for a more detailed introduction to the field.
@@ 2.1 Motion planning in dynamic environments
The motion planning problem is frequently solved with the conjunction of a global planner, which computes a static path considering a continuous and static map, and a local planner that follows it accounting for the dynamics of the robot and obstacles that were not considered by the global planner. Traditional local planners like the Dynamic Window Approach (DWA) [5] do not consider that obstacles may move in time, thus leading to collisions and suboptimal trajectories, as seen in some works as the benchmark proposed in [6]. First approaches for dynamic environments use attractive forces to lead the robot to its goal and repulsive forces to drive it away from obstacles, such as artificial potential fields [7] or Social Force [8]. They were later improved with Time Elastic Bands (TEB) [9] or multiple extensions to Social Force [10, 11].
A big group of works are velocity-space based. The Velocity Obstacle (VO), introduced in [12], refers to the set of possible robot velocities that could lead it to a collision with an obstacle that has a certain velocity in the near future. Based on the VO concept, the Dynamic Object Velocity Space (DOVS) for differential-drive robots is defined in [1] as the robot velocity space that includes the safe and unsafe robot possible velocities for a time horizon derived from VO. In that work, a planner based on strategies is also defined, the S-DOVS. Another work [13] uses the DOVS to discard velocities that lead to collisions from the action space of a tables-based reinforcement learning algorithm. However, unlike our work, it does not use the information of the DOVS to interpret the environment and just filters the unsafe actions available. In addition, the results presented in their work are very similar to S-DOVS, which does not require training time. Other extensions of VO use reciprocal collision assumptions, such as Reciprocal Velocity Obstacles (RVO) [14] or Optimal Reciprocal Collision Avoidance (ORCA) [15], which leads to failures in scenarios where there are non-collaborative obstacles.
@@ 2.2 Deep reinforcement learning planners
Deep Reinforcement Learning (DRL) [20] is a method used to learn to estimate the optimal policy that optimizes the cumulative reward obtained in an episode with a deep neural network. DRL algorithms have proven state-of-the-art performance in several benchmarks, including Rainbow [21] for discrete action spaces, distributed reinforcement learning algorithms [22] or actor-critic methods [23].
Some works study the benefits of DRL in robot motion planning and the limitations of traditional planners in dynamic environments. Defining strategies for every situation that may be found in the real world or optimizing trajectories in very dynamic environments is intractable, so DRL emerges to efficiently solve the decision-making problem, which is complex and has many degrees of freedom. There are works [6, 24, 25] that implement and compare DRL algorithms with model-based ones, proving their performance.
There are two big groups of DRL planners: end-to-end and agent-based planners. The former use the raw sensor measurements directly as the DRL network input. These measurements may be sensor maps [26, 27, 28, 29], LiDAR laser scans [30, 31, 32, 33] or images [34, 35, 36] taken from robot’s onboard sensors. In spite of the efforts in designing realistic simulators [37, 25, 38], these methods still experience distribution-shift problems when the deployment environments differ from the training one [39, 40], and require the robot to have the same sensor configuration (for example, sensor position, range or resolution) both in training and in deployment, limiting the applicability of their policy. Agent-based planners extract an agent-level representation from sensor measurements and use it as the planner input. In this way, the robot does not require to have specific sensor settings and does not experience distribution shifts due to the scenario geometry and shape. Furthermore, it may learn to understand deeper agent-level behaviors by differentiating decision-making agents from other obstacles [41].
The first agent-based DRL approach proposed for motion planning [42] uses a fully connected network to select velocities regarding the position and velocity of the closest surrounding obstacles. LSTM [43] layers were included in [41, 44] to account for multiple obstacles. SARL [45] uses attention to model interactions among obstacles, achieving state-of-the-art performance. Recent approaches compare their results to it, and they usually focus on improving the structure of the network. For example, [46] uses a graph neural network or [47] an attention-based interaction graph. In a recent study, [48] we showed that the performance of DRL approaches may be improved using intrinsic rewards. The use of DRL algorithms in robot navigation has many important challenges, including the need of learning from a limited subset of possible scenarios, environmental and robot constraints or partial observability [49, 50]. The previous approaches do not face these problems. In our approach, instead of directly process the observations, we use the DOVS model, that represents the dynamism of the environment in the velocity space, as an input to the DRL algorithm, doing the decision-making process agnostic of the specific scenario.
@@
3 Problem formulation
We consider a differential-drive robot that shares the workspace $\mathcal{W}\subseteq\mathbb{R}^{2}$ with static and dynamic obstacles. The robot position at time $t$ is $\mathbf{p_{t}}=(x_{t},y_{t},\theta_{t})\in\mathcal{W}\times[-\pi,\pi)$ and its velocity $\mathbf{u}_{t}=(\omega_{t},v_{t})\in\mathcal{V}\subseteq\mathbb{R}^{2}$ (angular and linear velocity). The robot maximum linear and angular velocities are bounded: $\mathcal{V}=[-\omega_{max},\omega_{max}]\times[0,v_{max}]$. It has to reach a goal $\mathbf{g}=(x_{g},y_{g})\in\mathcal{W}$ in the minimum possible time, while avoiding static and dynamic obstacles in the environment. The robot is represented with a disk with radius $r_{robot}$, and the dynamic obstacles are assumed to be circular disks with radius $\mathbf{r}_{obs}\in\mathbb{R}^{M}$, where $M$ is the number of dynamic obstacles. For each obstacle $i$, we assume that the robot may estimate its current position and velocity using the sensor information. Let $\mathcal{R}_{t}(\mathbf{u})\in\mathcal{W}\times[-\pi,\pi)$ be the robot state at time $t$ if it drives with velocity $\mathbf{u}$, $\mathscr{O}_{i,t}\in\mathcal{W}$ the set of points occupied by obstacle $i$ at time $t$, $\mathscr{O}_{i,t}^{T_{h}}=[\mathscr{O}_{i,t},\dots,\mathscr{O}_{i,t+T_{h}}]$ for the time horizon $T_{h}$ and $\mathscr{O}_{\forall,t}^{T_{h}}=[\mathscr{O}_{1,t}^{T_{h}},\dots,\mathscr{O}_{M,t}^{T_{h}}]$ for the whole set of static and dynamic obstacles in the environment. We consider that the robot follows the unicycle model:
@@ 3 Problem formulation
The robot is also restricted by the low-level forward model that relates the wheel velocities to the actions $\mathbf{u}_{t}$ for a differential-drive robot:
@@
where $r_{w}$ is the wheel radius, $\omega_{r,t}$ and $\omega_{l,t}$ the angular velocity of the right and left wheels, respectively, and $L_{w}$ the distance between wheels. The relationship $\frac{v_{t}}{\omega_{t}}=\frac{2}{L_{w}}$ imposes two linear constraints regarding the maximum linear velocity a robot may take in a particular moment with respect to its angular velocity:
@@
We consider that the robot has a maximum linear acceleration of $a_{max}$. The change in the linear velocity of the robot is limited by its previous linear velocity and, due to Equation 2, its velocity limits:
@@
The restrictions of Equation 3 and Equation 4 are plotted in a graphical representation of the robot velocity space in Figure 2. The velocity in the next control period $\boldsymbol{u}_{t+1}$ can never be above the two big black lines that relate linear and angular velocity constraints, and can never be outside the green dynamic window that surround $\boldsymbol{u}_{t}$, which constraints linear and angular acceleration and their relationship.
Figure 2: Graphical representation of the DOVS model and the differential-drive robot restrictions. Constraints of Equation 3 are represented with two black lines and restrict maximum linear velocities regarding the angular velocity. Equation 4 are plotted as a green rhombus around $\boldsymbol{u}_{t}$, representing acceleration limits regarding a differential-drive robot. In this example, the robot velocity limits are $v_{max}=0.7$ m/s, $\omega_{max}=\pi$ rad/s and $a_{max}=0.3$ m/s². The dark (DOV) and white (FV) areas include unsafe and safe velocities derived using VO for a time horizon.
4 Methodology
In this section, we explain the concepts of the DOVS and its adaptation used in RUMOR, as well as the deep reinforcement learning solution proposed to solve the problem.
@@ 4.1 Dynamic Object Velocity Space (DOVS)
In this work, we use the DOVS [1] to model the information of the environment. It translates scenario information from the workspace $\mathcal{W}$ to the robot velocity space $\mathcal{V}$, by computing the velocities in $\mathcal{V}$ lead to a collision within a temporal horizon. This provides a robocentric, bounded, scalable and flexible model with the original information of the workspace.
The DOVS is modeled using possible robot velocities. As we consider a differential-drive robot, its velocities result in circular trajectories if they are kept constant. We denote $\tau_{j}\equiv\tau_{j}(\omega_{j},v_{j})$ the robot trajectory $\tau_{j}$ produced by $\boldsymbol{u}_{j}=(\omega_{j},v_{j})\in\mathcal{V}$ and $\mathcal{T}$ the collection of different circular trajectories that collectively cover $\mathcal{V}$. The set of colliding velocities with the obstacle $i$ are the trajectories that intersect with $i$ within a time horizon. They define Dynamic Object Velocity of $i$:
Definition 1.
The Dynamic Object Velocity (DOV) of obstacle $i$, for the set of trajectories $\mathcal{T}$ and the time horizon $T_{h}$ is defined as:
$\begin{split}DOV(\mathscr{O}_{i,t}^{T_{h}},\mathcal{T})=&\left\{\mathbf{u}_{j}=(\omega_{j},v_{j})\in\mathcal{V}\mid\exists\tau_{j}\in\mathcal{T},\;\right.\\ &\left.\exists t^{\prime}\in[t,t+T_{h}],\;\mathcal{R}_{t^{\prime}}(\mathbf{u}_{j})\cap\mathscr{O}_{i,t^{\prime}}\neq\emptyset\right\}\\ \end{split}$ (5)
The DOV of all obstacles in the environment is combined for the general definition:
Definition 2.
The Dynamic Object Velocity (DOV) of $M$ obstacles of the set of trajectories $\mathcal{T}$ and the time horizon $T_{h}$ is defined as:
$DOV(\mathscr{O}_{\forall,t}^{T_{h}},\mathcal{T})=\bigcup_{i=1}^{M}DOV(\mathscr{O}_{i,t}^{T_{h}},\mathcal{T})$ (6)
The Free Velocity (FV) is constructed as the complementary set of the DOV with the collision-free velocities:
Definition 3.
The Free Velocity (FV), for the set of trajectories $\mathcal{T}$ and the time horizon $T_{h}$ is defined as:
$\begin{split}FV(\mathscr{O}_{\forall,t}^{T_{h}},\mathcal{T})=&\left\{\mathbf{u}_{j}=(\omega_{j},v_{j})\in\mathcal{V}\mid\mathbf{u}_{j}\notin DOV(\mathscr{O}_{\forall,t},\mathcal{T})\right\}\\ \end{split}$ (7)
Finally, the DOVS is modeled as the union of the whole set of velocities that comprise the DOV and the FV, categorizing every velocity in $\mathcal{V}$:
Definition 4.
The Dynamic Object Velocity Space (DOVS), for the set of trajectories $\mathcal{T}$ and the time horizon $T_{h}$ is defined as:
$DOVS(\mathscr{O}_{\forall,t}^{T_{h}},\mathcal{T})=\left\{DOV(\mathscr{O}_{\forall,t}^{T_{h}},\mathcal{T})\cup FV(\mathscr{O}_{\forall,t}^{T_{h}},\mathcal{T})\right\}$ (8)
The DOVS is computed with the concept of the collision band, $\mathcal{B}_{i}$. The robot is reduced to a point and the obstacle is enlarged with the robot radius, to simplify computations while keeping the same collision area. The collision band, $\mathcal{B}_{i}$, is the area that obstacle $i$ will occupy by following its predicted trajectory. In this work, we assume the robot only knows obstacle $i$ current velocity, so its trajectory is predicted by considering that its velocity is kept constant. Thus, obstacle trajectories are straight lines or circular trajectories, depending on whether the obstacle has angular velocity or not. The DOVS is computed with the intersections of trajectories and $\mathcal{B}_{i}$. For each robot circular trajectory $\tau_{j}\in\mathcal{T}$, the intersection points between the two lines limiting $\mathcal{B}_{i}$ and $\tau_{j}$ are $P_{1,i,j}(x_{1,i,j},y_{1,i,j})$ and $P_{2,i,j}(x_{2,i,j},y_{2,i,j})$. They are graphically represented in Figure 3 (a) in a robocentric view for a few example trajectories.
@@
Figure 3: A robocentric view of $\mathcal{W}$ (a) and a graphical representation of the DOVS (b) of a scenario where a robot faces the obstacle $i$ that follows a linear trajectory. In (a), the robot center is represented in red, the trajectories $\tau_{j}\in\mathcal{T}$ sampled in blue lines, the obstacle augmented radius with a blue circle and the collision band $\mathcal{B}_{i}$ with green lines. The intersection points between $\tau_{j}$ and $\mathcal{B}_{i}$ are $P_{1,i,j}$ (right) and $P_{2,i,j}$ (left). In (b), the maximum velocities to pass after the obstacle are represented with black dots ($\boldsymbol{u}_{2,i,j}$) and the minimum velocities to pass before it with purple dots ($\boldsymbol{u}_{1,i,j}$). The DOV is represented in gray and the FV in white.
Different combinations of linear and angular velocities lead to the same trajectory, as long as the trajectory radius, $\frac{v}{\omega}$, is the same. Depending on the velocity, the robot will eventually be at the same time as obstacle $i$ in $\mathcal{B}_{i}$, colliding, or will not. The time $t_{2,i,j}$, which is the time any point of the obstacle $i$ takes to reach $P_{2,i,j}$, and $t_{1,i,j}$, which is the time it takes to stop being in contact to $P_{1,i,j}$, may be efficiently computed considering the velocity of $i$ [1]. We denote $\boldsymbol{u}_{k,i,j}$ the velocity that makes the robot follow trajectory $\boldsymbol{\tau}_{j}$ and reach $P_{k,i,j}$ at time $t_{k,i,j}$, for $k=1,2$. Colliding velocities are velocities $\boldsymbol{u}_{j}$ that follow $\boldsymbol{\tau}_{j}$ and $v_{2,i,j}\leq v_{j}\leq v_{1,i,j}$. Therefore, $v_{1,i,j}$ is the robot minimum linear velocity to pass before the obstacle passes without collision, and $v_{2,i,j}$ the maximum linear velocity to pass after it. They are represented in Figure 3(b), and they are computed as:
@@
The computation of $\boldsymbol{u}_{k,i,j}=(\omega_{k,i,j},v_{k,i,j})$ for $\mathcal{T}$ and all the obstacles sets the limits of the DOV. The union of the DOV and the free robot velocity space is the DOVS, represented in Figure 3.
Remark 1.
The DOVS is used for representing information of the scenario accessible by the robot in $\mathcal{V}$. This representation is therefore valid even though the future trajectories of obstacles are unknown, and the obstacles change their velocities in time. The DOVS is recomputed every control period with the updated information, which is then used for the decision making.
The differential-drive velocity constrains of Equations 3 and 4 are directly constructed in the DOVS, as shown in Figure 2. This combination is used to learn how to navigate safely, as explained in the following sections.
4.2 Reinforcement learning setup